OpenStudy (love_333):

Medal + fan: The volume of a sphere is 4,000pi m^3. What is the surface area of the sphere to the nearest square meter? A 2,614m^2 B 3,836m^2 C 1,307m^2 D 764m^2

OpenStudy (love_333):

@Erak

OpenStudy (erak):

So find the radius when given volume of a sphere. V = (4/3)pi(r)^3

OpenStudy (love_333):

How exactly do I set this one up?

OpenStudy (erak):

So you're given volume V = (4/3)pi(r)^3 4000pi = (4/3)pi(r)^3 I would just cross out pi, as it's on both sides, so solve this for r 4000 = (4/3)(r)^3

OpenStudy (love_333):

When I do it I get 9.847

OpenStudy (love_333):

I feel like I'm not doing it correctly...

OpenStudy (erak):

so divide both sides by (4/3), which basically is 4000*3/4, then cube root it

OpenStudy (love_333):

I got 3,000 and then 14.422

OpenStudy (erak):

yeah

OpenStudy (erak):

So that's your radius, put it into the SA formula

OpenStudy (love_333):

I got 2,461.76?

OpenStudy (love_333):

What if thats not one of my choices? @Erak

OpenStudy (erak):

So that's probably cause you rounded somewhere. Logically, it's close enough to A that you could guess that based on what you got

OpenStudy (love_333):

Okay, that makes sense but i'm not sure what I did.

OpenStudy (erak):

So does your calculator have an "Ans" function? Go back to getting your radius of 14.422, then do 4pi*ans^2